OpenStudy (anonymous):

which is equal to 2 sqrt -27? A. 6 sqrt 3 B. -6 sqrt 3 C. -6i sqrt 3 D. 6i sqrt 3

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (anonymous):

i'm usually pretty good with math, and don't normally ask questions, but i absolutely don't understand the concept of imaginary numbers. i take an online math class, so the explanation is pretty vague... so no, no ideas at all :(

OpenStudy (skullpatrol):

By definition i^2 = -1.

OpenStudy (anonymous):

okay, so how does that apply to the radical?

OpenStudy (anonymous):

like how would I solve it?

OpenStudy (whpalmer4):

\[i^2=-1\]\[i=\sqrt{-1}\] \[\sqrt{-27}=\sqrt{27*(-1)}=\sqrt{9*(-1)*3}=\sqrt{9*i^2*3}=3i\sqrt{3}\]

OpenStudy (anonymous):

wouldn't it be 6i sqrt 3?

OpenStudy (whpalmer4):

I was only doing sqrt(-27) to illustrate how to evaluate the radical, your problem is 2 sqrt(-27). Had to leave some of the work for you :-)

OpenStudy (anonymous):

i figured it out after :) thank you

OpenStudy (whpalmer4):

Trivia factoid for the night: electrical engineers make extensive use of imaginary numbers, but they don't use i — they use j instead :-) Same value. i gets used for current.

OpenStudy (whpalmer4):

More useful bit of trivia: \[e^{i\pi}+1=0\]How cool is that?